OpenStudy (anonymous):

Solve: -20 - 3a = - 5(a - 2)

OpenStudy (bahrom7893):

simplify the right side: -20-3a = -5a+10 move all as to one side and all numbers to another: -3a+5a=10+20 simplify further: 2a=30 divide both sides by 2: a=15

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

firstly distribute that -5 to (a-2) so -5(a-2)=-5a+10. Now you have -20-3a=-5a+10. your goal is to get the a in one side and the digits on the other. bahrom7893 has it right on

OpenStudy (anonymous):

Thank you both, very much.

OpenStudy (anonymous):

if you ever need any help with these type of problems, let me know some how i think O.o